# Sub-GeV millicharge dark matter from the $U(1)_{X}$ hidden sector

### v2

@@ 1 Introduction
The connection between kinetic mixing and the millicharge can be traced back to [28, 29]. The kinetic mixing between a massless $U(1)_{a}$ and a massless $U(1)_{b}$ can generate a milli-$U(1)_{a}$ charge for a $U(1)_{b}$-charged fermion, proportional to the kinetic mixing parameter. In a massive amount of literature, the generation of millicharge particle boils down to a simplified effective model that an extra $U(1)_{X}$ mixed with the photon field through gauge kinetic terms, i.e., via the effective operator $\mathcal{L}\sim\frac{\bm{\delta}}{2}F_{\mu\nu}^{\rm em}F_{X}^{\mu\nu}$. After the mixing, fermions charged under the extra $U(1)_{X}$ will carry a millicharge. However, as we will review in Section 2.2, when one of the $U(1)$’s gets massive, say $U(1)_{b}$, the kinetic mixing with the massless $U(1)_{a}$ can no longer generate a milli-$U(1)_{a}$ charge for a $U(1)_{b}$-charged fermion, no matter how tiny the massive $U(1)_{b}$ is.
In the full electroweak theory, the photon field is a linear combination of the hypercharge $B$ and the neutral $SU(2)$ gauge fields $A_{3}$. Thus the kinetic mixing between the extra $U(1)$ and the hypercharge gauge fields should be treated simultaneously with the $\{B,A_{3}\}$ electroweak mass mixing matrix. It has been shown that pure kinetic mixing between an extra $U(1)$ and the hypercharge gauge field cannot generate a millicharge [30]. More precisely, for a general discussion of the Standard Model (SM) extended by a $U(1)_{X}$ hidden sector, the millicharge carried by the $U(1)_{X}$ dark fermions can be obtained from: (1) The dark particle carries a tiny amount of hypercharge as a prior. (2) A kinetic mixing between a massless $U(1)$ and the hypercharge gauge field, and the generated millicharge is proportional to the kinetic mixing parameter. (3) The mass mixing between a massive $U(1)$ with the hypercharge gauge field [31, 30]. In the last case the kinetic mixing does not play any role in generating the millicharge, and the millicharge generated is proportional to the mass mixing parameter. In this paper we will review all these possibilities and will focus on the third case, exploring details in a realistic model involving viable freeze-in millicharge dark matter from a $U(1)_{X}$ hidden sector. For a realistic model, it is essential to ensure that in the final mass eigenbasis, the photon is precisely massless.
@@ 2 The $U(1)$ extension of the SM and the generation of millicharge particle
We outline our primary conclusion at the beginning that the millicharge particle can be generated by mixing only in two ways: (1) kinetic mixing between two massless $U(1)$’s [28], and (2) mass mixing between two massive $U(1)$’s (though in the mass eigenbasis there can still be a massless $U(1)$[31, 30]. We emphasize here that except for the two massless $U(1)$ case, the kinetic mixing does not play any role in generating the millicharge carried by the dark particle. For the case of two massive $U(1)$’s, if kinetic mixing and mass mixing are both evoked, the millicharge carried by the fermion will only depend on the mass mixing.
@@
In the following we would like to summarize some of the known results regarding the generation of millicharge particles, and we wish to clarify several important issues as well as some misunderstandings regarding the generation of the millicharge with kinetic and/or mass mixing. We will first review the mixing effect for two $U(1)$’s and then discuss the mixing effect of an extra $U(1)_{X}$ with the full electroweak theory.
2.1 Kinetic mixing between two massless $U(1)$’s
The simplest case of kinetic mixing arises from two massless $U(1)$’s [28, 29]. From an effective theory approach, at low energies one can consider the case that an extra $U(1)_{X}$ mixed with the photon field. In the gauge eigenbasis, the kinetic part and interaction part of the Lagrangian read
@@ 2.1 Kinetic mixing between two massless $U(1)$’s
One can see now the dark current can couple to photon with strength $s_{\bm{\delta}}g_{X}$, corresponding to a millicharge $s_{\bm{\delta}}g_{X}/e$ carried by $\chi$ (we assume $\chi$ carries the $U(1)_{X}$ charge $+1$). However, the transformation Eq. (9) is arbitrary, one can reverse $A,C$ in Eq. (10) as
@@
and one can see now the SM fermions carry milli-$U(1)_{X}$ charges $s_{\bm{\delta}}eQ_{i}/g_{X}$, where $Q_{i}$ denotes the electric charges carried by SM fermions in the unit of $e$, though this transformation is phenomenologically not interesting.
@@ 2.2 Kinetic mixing between a massless $U(1)$ and a massive $U(1)$
A more realistic case is kinetic mixing between the photon field with a massive $U(1)$. In this case the Lagrangian reads
@@
Now the only possible way of eliminating the kinetic mixing term is given by the following transformation
$\left(\begin{array}[]{c}C\\ A\end{array}\right)=\left(\begin{array}[]{cc}c_{\bm{\delta}}&0\\ -s_{\bm{\delta}}&1\end{array}\right)\left(\begin{array}[]{c}A_{X}\\ A_{\gamma}\end{array}\right)\,,$ (21)
similar to Eq. (12), to maintain the physical photon field exactly massless. In other words, the nonzero mass of the massive $U(1)$ field fixes the angle $\alpha$ to be $\pi/2$ and thus fixes the rotation matrix $R$ in Eq. (14). The above transformation gives rise to the interaction in the physical eigenbasis as
and one can see in this case, the dark particles carry exactly zero electric charge.
Hence, it is not appropriate to consider a kinetic mixing between the photon and a massive dark photon, and still claiming the dark particles from the dark $U(1)$ sector can achieve a millicharge proportional to the kinetic mixing parameter $\bm{\delta}$. More precisely, the case of kinetic mixing between two massless $U(1)$’s, is distinct from the situation where a massless $U(1)$ kinetically mixed with a massive $U(1)$, even when the latter has an extremely small but nonzero mass $M$. One cannot take the $M\to 0$ limit in this case and use directly the result from the case of kinetic mixing between two massless $U(1)$’s. In summary, for a dark photon with a non-negligible mass (for example, a $\mathcal{O}$(keV–MeV) dark photon as commonly studied), even though this mass is much smaller than the experimental scale one considered, the kinetic mixing between a massless $U(1)$ and a massive $U(1)$ does not generate a millicharge carried by the dark particle.
As long as one of the $U(1)$’s is massive, the only possible way of generating the millicharge through mixing is from the mass mixing effect, and we will review this mechanism in the next subsection.
2.3 Kinetic mixing and mass mixing between two massive $U(1)$’s
The Stueckelberg mass mixing effect of two massive $U(1)$ is discussed in [30]. The full Lagrangian reads
@@ 2.3 Kinetic mixing and mass mixing between two massive $U(1)$’s
$\mathcal{K}=\left(\begin{array}[]{cc}1&\bm{\delta}\\ \bm{\delta}&1\end{array}\right)\,,\qquad M_{{\rm St}}^{2}=\left(\begin{array}[]{cc}M_{1}^{2}&M_{1}M_{2}\\ M_{1}M_{2}&M_{2}^{2}\end{array}\right)=M_{1}^{2}\left(\begin{array}[]{cc}1&\epsilon\\ \epsilon&\epsilon^{2}\end{array}\right)\,,$ (26)
@@
$V^{T}M_{{\rm St}}^{2}V\to(KV^{\prime})^{T}M_{{\rm St}}^{2}(KV^{\prime})=V^{\prime T}K^{T}M_{{\rm St}}^{2}KV^{\prime}\,.$ (29)
Performing one further orthogonal transformation $O$ to diagonalize the matrix $K^{T}M_{{\rm St}}^{2}K$ such that
$O^{T}K^{T}M_{{\rm St}}^{2}KO=D^{2}\,,$ (30)
@@
$V^{\prime T}K^{T}M_{{\rm St}}^{2}KV^{\prime}=V^{\prime T}O(O^{T}K^{T}M_{{\rm St}}^{2}KO)O^{T}V^{\prime}=(V^{\prime T}O)D^{2}(O^{T}V^{\prime})\equiv E^{T}D^{2}E\,,$ (31)
@@
2.4 Kinetic mixing and mass mixing in the full electroweak theory
@@ 2.4.1 $U(1)_{X}$ gauge boson achieves a mass via the Stueckelberg mechanism
$a_{f}$ $=\left(g_{2}\mathcal{R}_{33}-g_{Y}\mathcal{R}_{23}\right)T_{i}^{3}\,,$ (41)
@@
$a_{f}^{\prime}$ $=\left(g_{2}\mathcal{R}_{31}-g_{Y}\mathcal{R}_{21}\right)T_{i}^{3}\,.$ (43)
@@ 3 Coupled Boltzmann equations with multiple temperatures
In Fig. 2, we graphically illustrate the model we discuss. Initially there is negligible amount of matter in the hidden sector. The dark fermion $\chi$ and the dark photon $\gamma^{\prime}$ are produced through the freeze-in production of the SM particles. Once the matter inside the hidden sector accumulated, hidden sector interactions start to take place, such as $\chi\bar{\chi}\leftrightarrow\gamma^{\prime}\gamma^{\prime}$ which will change the number densities of $\chi$ and $\gamma^{\prime}$. The hidden sector interactions occur simultaneously with the freeze-in processes. The hidden sector possesses a different temperature $T_{h}$, which is related to the visible sector temperature $T_{v}$ (the temperature of the observed Universe, and we will omit the subscript ${}_{v}$ from here), with the function $\eta(T_{h})$. When the dark photon mass is greater than twice of the electron mass, the dark photon will eventually decay to (mostly) electron-positron pairs. When the dark photon mass is less than twice of electron mass, it can only decay into two neutrinos and to three photons and must be made stable through the age of the Universe, see discussions in Section 2.4.1.
@@ 3.1 Derivation of the evolution of hidden sector particles
3.2 Coupled Boltzmann equations from two sectors
@@
where the upper subscript is the visible sector temperature $T$ or hidden sector temperature $T_{h}$ that the corresponding process is depending on. The freeze-in processes are marked by green color to be distinguishable to various hidden sector processes. Later we will use $T_{h}$ as the reference temperature and replace $T$ by $T_{h}\eta$. In the above equations, $i$ and $\bar{i}$ represent SM fermions and anti-fermions, $\gamma^{*}$ denotes the plasmon which expresses photon with effective in-medium thermal mass in the early universe, c.f., Appendix B for a brief review. All the terms in the above equations with subscript ${}^{{\rm eq}}$ indicate that we are calculating the corresponding process reversely for computational simplicity. In such calculations, we use the equilibrium values of comoving number densities (depending on the visible sector temperature $T$)
@@
Figure 3: A display of production channels for the dark fermion $\chi$, including the SM fermion annihilation $i\bar{i}\to\chi\bar{\chi}$ through $s$ and $t$ channels, as well as the plasmon decay $\gamma^{*}\to\chi\bar{\chi}$.
@@ 4.1 $U(1)_{X}$ sub-GeV dark matter via freeze-in
In this case hidden sector particles interchange mostly through the three-point interactions $\gamma^{\prime}\leftrightarrow\chi\bar{\chi}$. Because of the rather strong interactions among the hidden sector, the dark photon will soon decay to $\chi\bar{\chi}$ when the temperature drops down. Four-point interactions $\chi\bar{\chi}\leftrightarrow\gamma^{\prime}\gamma^{\prime}$ can be neglected in this case. In addition, the decay channel of dark photon to SM fermions can also be neglected since the decay of $\gamma^{\prime}\to\chi\bar{\chi}$ is so much stronger.
@@ 5 Conclusion
Inappropriate treatments and misunderstandings include:
@@
    Even from the effective field theory point of view, the kinetic mixing between the massless photon field and an extra massive $U(1)$ cannot generate a millicharge, no matter how tiny the extra $U(1)$ mass is [30], summarised in Section 2.2. Taking the mass of the massive $U(1)$ to be extremely small (but still nonzero), cannot transit to the case of kinetic mixing between two massless $U(1)$’s, which will generate a millicharge proportional to the kinetic mixing parameter.
@@
    The kinetic mixing does not play any role in generating a millicharge as long as one of the two $U(1)$’s gets massive. In this case only the mass mixing, e.g., Eq. (24) can generate a millicharge.

### v3

@@ 1 Introduction
The connection between kinetic mixing and the millicharge can be traced back to [28, 29]. The kinetic mixing between a massless $U(1)_{a}$ and a massless $U(1)_{b}$ can generate a milli-$U(1)_{a}$ charge for a $U(1)_{b}$-charged fermion, proportional to the kinetic mixing parameter. In a massive amount of literature, the generation of millicharge particle boils down to a simplified effective model that an extra $U(1)_{X}$ mixed with the photon field through gauge kinetic terms, i.e., via the effective operator $\mathcal{L}\sim\frac{\bm{\delta}}{2}F_{\mu\nu}^{\rm em}F_{X}^{\mu\nu}$. After the mixing, fermions charged under the extra $U(1)_{X}$ will carry a millicharge. However, as we will review in Section 2.2, when one of the $U(1)$ fields gets massive, say $U(1)_{b}$, the kinetic mixing with the massless $U(1)_{a}$ can no longer generate a milli-$U(1)_{a}$ charge for a $U(1)_{b}$-charged fermion, no matter how tiny the massive $U(1)_{b}$ is.
In the full electroweak theory, the photon field is a linear combination of the hypercharge $B$ and the neutral $SU(2)$ gauge fields $A_{3}$. Thus the kinetic mixing between the extra $U(1)$ and the hypercharge gauge fields should be treated simultaneously with the $\{B,A_{3}\}$ electroweak mass mixing matrix [30]. It has been shown that pure kinetic mixing between an extra $U(1)$ and the hypercharge gauge field cannot generate a millicharge [31]. More precisely, for a general discussion of the Standard Model (SM) extended by a $U(1)_{X}$ hidden sector, the millicharge carried by the $U(1)_{X}$ dark fermions can be obtained from: (1) The dark particle carries a tiny amount of hypercharge as a prior. (2) A kinetic mixing between a massless $U(1)$ and the hypercharge gauge field, and the generated millicharge is proportional to the kinetic mixing parameter. (3) The mass mixing between a massive $U(1)$ with the hypercharge gauge field [32, 31]. In the last case the kinetic mixing does not play any role in generating the millicharge, and the millicharge generated is proportional to the mass mixing parameter. In this paper we will review all these possibilities and will focus on the third case, exploring details in a realistic model involving viable freeze-in millicharge dark matter from a $U(1)_{X}$ hidden sector. For a realistic model, it is essential to ensure that in the final mass eigenbasis, the photon is precisely massless.
@@ 2 The $U(1)$ extension of the SM and the generation of millicharge particle
We outline our primary conclusion at the beginning that the millicharge particle can be generated by mixing only in two ways: (1) kinetic mixing between two massless $U(1)$ fields [28], and (2) mass mixing between two massive $U(1)$ fields (though in the mass eigenbasis there can still be a massless $U(1)$[32, 31]. We emphasize here that except for the two massless $U(1)$ case, the kinetic mixing does not play any role in generating the millicharge carried by the dark particle. For the case of two massive $U(1)$ fields, if kinetic mixing and mass mixing are both evoked, the millicharge carried by the fermion will only depend on the mass mixing.
@@
In the following we would like to summarize some of the known results regarding the generation of millicharge particles, and we wish to clarify several important issues as well as some misunderstandings regarding the generation of the millicharge with kinetic and/or mass mixing. We will first review the mixing effect for two $U(1)$ fields and then discuss the mixing effect of an extra $U(1)_{X}$ with the full electroweak theory.
2.1 Kinetic mixing between two massless $U(1)$ fields
The simplest case of kinetic mixing arises from two massless $U(1)$ fields [28, 29]. From an effective theory approach, at low energies one can consider the case that an extra $U(1)_{X}$ mixed with the photon field. In the gauge eigenbasis, the kinetic part and interaction part of the Lagrangian read
@@ 2.1 Kinetic mixing between two massless $U(1)$ fields
One can see now the dark current can couple to photon with strength $s_{\bm{\delta}}g_{X}$, corresponding to a millicharge $s_{\bm{\delta}}g_{X}/e$ carried by $\chi$ (we assume $\chi$ carries the $U(1)_{X}$ charge $+1$). However, the non-unitary transformation Eq. (9) is not symmetric with respect to the fields $A,C$. Thus instead of the transformation made in Eq. (10), reversing the order of $A,C$ one has
@@
corresponding to the case where dark particles carry zero electric charge (no millicharge), while SM fermions carry milli-$U(1)_{X}$ charges $s_{\bm{\delta}}eQ_{i}/g_{X}$ ($Q_{i}$ denotes the electric charges carried by SM fermions in the unit of $e$).
Thus for the kinetic mixing between two massless $U(1)$ fields, there can be two distinct outcomes: (1) Dark particles carry millicharge while SM particles carry no $U(1)_{X}$ charge. (2) Dark particles carry no millicharge while SM particles carry milli-$U(1)_{X}$ charge. This characteristic arises due to the asymmetry in the non-unitary transformation. These two possibilities are both hypothetical and will be confirmed or falsified through experimentation. However, to generate a millicharge carried by dark particles, the transformation in Eq. (10) must be applied.
@@ 2.2 Kinetic mixing between a massless $U(1)$ and a massive $U(1)$
A more realistic case is kinetic mixing between the photon field with a massive $U(1)_{X}$. In this case the Lagrangian reads
@@
To maintain the physical photon field exactly massless, the only possible way of eliminating the kinetic mixing term is given by the following transformation
The above transformation gives rise to the interaction in the physical eigenbasis as
$\mathcal{L}_{{\rm int}}=g_{X}c_{\bm{\delta}}A_{X}J_{{\rm d}}+e(-s_{\bm{\delta}}A_{X}+A_{\gamma})J_{{\rm em}}\,.$ (22)
It is clear that dark particles carry precisely zero electric charge in this case.
As mentioned in the previous subsection, the non-unitary transformation Eq. (9) is not symmetric with respect to the fields $A,C$. For the kinetic mixing between two massless $U(1)$ fields, there are two ways of taking the transformation Eqs. (10) and (12), one leads to dark particles carry the millicharge and the other leads to no millicharge. However, the emergence of the mass term for the massive $U(1)$ determines the transformation to be Eq. (21), ensuring that the physical photon remains precisely massless. The transformation Eq. (21) in the current case is identical to the transformation in Eq. (12) for two massless $U(1)$ fields, which leads to no millicharge for dark particles.
Hence, it is not appropriate to consider a kinetic mixing between the photon and a massive dark photon, and still claiming the dark particles from the dark $U(1)$ sector can achieve a millicharge proportional to the kinetic mixing parameter $\bm{\delta}$. The generation of millicharge from the case of kinetic mixing between two massless $U(1)$ fields is distinct from the situation where a massless $U(1)$ kinetically mixed with a massive $U(1)_{X}$, even when the latter has an extremely small but nonzero mass $M$. One cannot take the $M\to 0$ limit in this case and use directly the result from the case of kinetic mixing between two massless $U(1)$ fields in Eq. (11), where the millicharge carried by dark particles is generated, since these two cases used different transformations shown in Eq. (21) and Eq. (10). The $M\to 0$ limit brings us to the transformation in Eq. (12) for two massless $U(1)$ fields, resulting in no millicharge being produced for the dark particles while SM particles carry milli-$U(1)_{X}$ charge.
In summary, for a dark photon with a non-negligible mass (for example, a $\mathcal{O}$(keV–MeV) dark photon as commonly studied), even though this mass is much smaller than the experimental scale one considered, the kinetic mixing between a massless $U(1)$ and a massive $U(1)$ does not generate a millicharge carried by the dark particle.
As long as one of the $U(1)$ fields is massive, the only possible way of generating the millicharge through mixing is from the mass mixing effect, and we will review this mechanism in the next subsection.
2.3 Kinetic and mass mixing between two massive $U(1)$ fields
The mass mixing effect of two massive $U(1)$ is discussed in [30, 31]. The full Lagrangian reads
@@ 2.3 Kinetic and mass mixing between two massive $U(1)$ fields
$\mathcal{K}=\left(\begin{array}[]{cc}1&\bm{\delta}\\ \bm{\delta}&1\end{array}\right)\,,\qquad M^{2}=\left(\begin{array}[]{cc}M_{1}^{2}&M_{1}M_{2}\\ M_{1}M_{2}&M_{2}^{2}\end{array}\right)=M_{1}^{2}\left(\begin{array}[]{cc}1&\epsilon\\ \epsilon&\epsilon^{2}\end{array}\right)\,,$ (26)
@@
$V^{T}M^{2}V\to(KV^{\prime})^{T}M^{2}(KV^{\prime})=V^{\prime T}K^{T}M^{2}KV^{\prime}\,.$ (29)
Performing one further orthogonal transformation $O$ to diagonalize the matrix $K^{T}M^{2}K$ such that
$O^{T}K^{T}M^{2}KO=D^{2}\,,$ (30)
@@
$V^{\prime T}K^{T}M^{2}KV^{\prime}=V^{\prime T}O(O^{T}K^{T}M^{2}KO)O^{T}V^{\prime}=(V^{\prime T}O)D^{2}(O^{T}V^{\prime})\equiv E^{T}D^{2}E\,,$ (31)
@@
2.4 Kinetic and mass mixing in the full electroweak theory
@@ 2.4.1 $U(1)_{X}$ gauge boson achieves a mass via the Stueckelberg mechanism
$a_{i}$ $=\left(g_{2}\mathcal{R}_{33}-g_{Y}\mathcal{R}_{23}\right)T_{i}^{3}\,,$ (41)
@@
$a_{i}^{\prime}$ $=\left(g_{2}\mathcal{R}_{31}-g_{Y}\mathcal{R}_{21}\right)T_{i}^{3}\,.$ (43)
@@ 3 Coupled Boltzmann equations with multiple temperatures
In Fig. 2, we graphically illustrate the model we discuss. Initially there is negligible amount of matter in the hidden sector. The dark fermion $\chi$ and the dark photon $\gamma^{\prime}$ are produced through the freeze-in production of the SM particles. Once the matter inside the hidden sector accumulated, hidden sector interactions start to take place, such as $\chi\bar{\chi}\leftrightarrow\gamma^{\prime}\gamma^{\prime}$ which will change the number densities of $\chi$ and $\gamma^{\prime}$. The hidden sector interactions occur simultaneously with the freeze-in processes. The hidden sector possesses a different temperature $T_{h}$, which is related to the visible sector temperature $T_{v}$ (the temperature of the observed Universe, and we will omit the subscript v from here), with the function $\eta(T_{h})$. When the dark photon mass is greater than twice of the electron mass, the dark photon will eventually decay to (mostly) electron-positron pairs. When the dark photon mass is less than twice of electron mass, it can only decay into two neutrinos and to three photons and must be made stable through the age of the Universe, see discussions in Section 2.4.1.
@@ 3.1 Derivation of the evolution of hidden sector particles
3.2 Coupled Boltzmann equations for two sectors
@@
where the upper subscript is the visible sector temperature $T$ or hidden sector temperature $T_{h}$ that the corresponding process is depending on. The freeze-in processes are marked by green color to be distinguishable to various hidden sector processes. Later we will use $T_{h}$ as the reference temperature and replace $T$ by $T_{h}\eta$. In the above equations, $i$ and $\bar{i}$ represent SM fermions and anti-fermions, $\gamma^{*}$ denotes the plasmon which expresses photon with effective in-medium thermal mass in the early universe, c.f., Appendix B for a brief review. All the terms in the above equations with subscript eq indicate that we are calculating the corresponding process reversely for computational simplicity. In such calculations, we use the equilibrium values of comoving number densities (depending on the visible sector temperature $T$)
@@
Figure 3: A display of production channels for the dark fermion $\chi$, including SM fermions annihilation $i\bar{i}\to\chi\bar{\chi}$ through $s$ channel, as well as the plasmon decay $\gamma^{*}\to\chi\bar{\chi}$.
@@ 4.1 $U(1)_{X}$ sub-GeV dark matter via freeze-in
In this case hidden sector particles interchange mostly through the three-point interactions $\gamma^{\prime}\leftrightarrow\chi\bar{\chi}$. Because of the rather strong interactions among the hidden sector, the dark photon will soon decay to $\chi\bar{\chi}$ when the temperature drops down. Four-point interactions $\chi\bar{\chi}\leftrightarrow\gamma^{\prime}\gamma^{\prime}$ can be neglected in this case. In addition, the decay channels of dark photon to SM fermions can also be neglected since the decay of $\gamma^{\prime}\to\chi\bar{\chi}$ is so much stronger.
@@ 5 Conclusion
We clarify some inappropriate treatments and misunderstandings that are presented in the literature:
@@
    Even from the effective field theory point of view, the kinetic mixing between the massless photon field and an extra massive $U(1)$ cannot generate a millicharge, no matter how tiny the extra $U(1)$ mass is [31], summarised in Section 2.2. Taking the mass of the massive $U(1)$ to be extremely small (but still nonzero), cannot transit to the case of kinetic mixing between two massless $U(1)$ fields, which will generate a millicharge proportional to the kinetic mixing parameter.
@@
    The kinetic mixing does not play any role in generating a millicharge as long as one of the two $U(1)$ fields gets massive. In this case only the mass mixing, e.g., Eq. (24) can generate a millicharge.